# Extending Nonlocal Kinetic Energy Density Functionals to Isolated Systems via a Density-Functional-Dependent Kernel

### v1

@@ Abstract
Introduction—Orbital-free density functional theory (OFDFT) [1, 2] circumvents the $O(N^{3})$ ($N$ is the electron number) computational bottleneck of Kohn-Sham DFT (KSDFT) [3, 4] by eliminating explicit orbital dependence, achieving near-linear scaling efficiency. The method has demonstrated promising applications in diverse systems including liquid metal surface structures [5], lightweight alloy design [6], and warm dense matter simulations [7, 8, 9]. However, current OFDFT still encounters enormous challenges in approximating the non-interacting kinetic energy $T_{s}[\rho]$ solely through the electron density $\rho(\mathbf{r})$ — a fundamental conundrum posed by the Hohenberg-Kohn theorem [3, 10]. Current kinetic energy density functionals (KEDFs) either sacrifice accuracy for universality or fail to maintain numerical stability in finite systems, severely restricting OFDFT’s applicability beyond bulk materials.
Semilocal KEDFs, constructed from the electron density $\rho(\mathbf{r})$, its dimensionless gradient $s=|\nabla\rho|/(2k_{\mathrm{F}}\rho)$ and Laplacian $q=\nabla^{2}\rho/(4k_{\mathrm{F}}^{2}\rho)$ (where $k_{\mathrm{F}}=(3\pi^{2}\rho)^{1/3}$), maintain universality across bulk and isolated systems through their locally defined variables. [11, 12, 13, 14, 15, 16, 17] Their neglect of nonlocal information, however, leads to systematic errors in key electronic properties: atomic shell structures vanish [1], charge density cusps at nuclei are unphysical unless additional correction introduced [18, 19].
Nonlocal KEDFs enhance accuracy by incorporating nonlocal information through convolution kernels $w(\mathbf{r},\mathbf{r}^{\prime})$. The Wang-Teter-like (WT-like) functionals, including WT [20], Smargiassi-Madden (SM) [21], and Perrot [22], employ a density-independent kernel $w(k_{\mathrm{F}}^{0},|\mathbf{r}-\mathbf{r}^{\prime}|)$, parametrized by $k_{\mathrm{F}}^{0}=(3\pi^{2}\rho_{0})^{1/3}$. Here, $\rho_{0}$ is a parameter conventionally set as the average charge density $\rho_{\mathrm{avg}}$. The WT-like KEDFs generally decompose the non-interacting kinetic energy $T_{s}$ as
@@
Distinct physical regimes motivate specific parameterizations: The original WT KEDF ($\alpha$=$\beta$=$\frac{5}{6}$) targets weakly varying uniform electron gas (UEG), while the Perrot KEDF ($\alpha$=$\beta$=1) was constructed for thin electron gas, and the SM KEDF ($\alpha$=$\beta$=$\frac{1}{2}$) is derived from low-$q$ limit. Despite achieving $O(N\ln N)$ scaling and improved accuracy in bulk systems, these functionals catastrophically failure in isolated systems due to the Blanc-Cancès (BC) instability [23], manifested as total energy unboundedness from below.
Our analysis (see Supplementary Information (SI)) traces the BC instability to the ill-defined nature of $\rho_{\mathrm{avg}}$ as a rigid spatial average that remains fixed under density scaling $\rho_{\sigma}(\mathbf{r})=\sigma^{3}\rho_{1}(\sigma\mathbf{r})$ in isolated systems. This unphysical constant simultaneously violates the exact scaling law $T_{s}[\rho_{\sigma}]=\sigma^{2}T_{s}[\rho_{1}]$ [24] and generates negative Pauli energy $T_{\theta}$ through its mismatch with the physical density $\rho(\mathbf{r})$, contradicting its fundamental positivity requirement [25]. The BC instability therefore stems from this dual failure of $\rho_{\mathrm{avg}}$, which both breaks scaling invariance and misrepresents the true density, collectively rendering the functional unbounded. This understanding addresses a longstanding challenge in nonlocal KEDF development.
@@
With the BC instability resolved, we benchmark the ext-WT KEDF against semilocal functionals, including the second-order gradient expansion (GE2) [14] and Luo-Karasiev-Trickey (LKT) KEDFs [16], alongside the nonlocal WT KEDF, and KSDFT for single-atom systems of 56 elements. All calculations were performed within the ABACUS 3.8.0 packages [32, 33] with Perdew-Burke-Ernzerhof (PBE) [34] exchange-correclation functional, employing three distinct potentials: bare Coulomb potential for H and helium (He), widely-used bulk-derived local pseudopotentials (BLPS) [35] covering nine elements, and recently proposed high-quality local pseudopotentials (HQLPS) [36] covering all simple and transition metals.

### v2

@@ Abstract
Introduction—Orbital-free density functional theory (OFDFT) Wang and Carter (2002); Mi et al. (2023) circumvents the $O(N^{3})$ ($N$ is the electron number) computational bottleneck of Kohn-Sham DFT (KSDFT) Hohenberg and Kohn (1964); Kohn and Sham (1965) by eliminating explicit orbital dependence, achieving near-linear scaling efficiency. The method has demonstrated promising applications in diverse systems including liquid metal surface structures González, González, and Stott (2004), lightweight alloy design Shin and Carter (2014), and warm dense matter simulations White et al. (2013); Sjostrom and Daligault (2014); Ding et al. (2018). However, current OFDFT still encounters enormous challenges in approximating the non-interacting kinetic energy $T_{s}[\rho]$ solely through the electron density $\rho(\mathbf{r})$ — a fundamental conundrum posed by the Hohenberg-Kohn theorem Hohenberg and Kohn (1964); Ruzsinszky and Perdew (2011). Current kinetic energy density functionals (KEDFs) either sacrifice accuracy for universality or fail to maintain numerical stability in finite systems, severely restricting OFDFT’s applicability beyond bulk materials.
Semilocal KEDFs, constructed from the electron density $\rho(\mathbf{r})$, its dimensionless gradient $s=|\nabla\rho|/(2k_{\mathrm{F}}\rho)$ and Laplacian $q=\nabla^{2}\rho/(4k_{\mathrm{F}}^{2}\rho)$ (where $k_{\mathrm{F}}=(3\pi^{2}\rho)^{1/3}$), maintain universality across bulk and isolated systems through their locally defined variables. Thomas (1927); Fermi (1927); Weizsäcker (1935); Kirzhnits (1957); Constantin et al. (2011); Luo, Karasiev, and Trickey (2018); Wang, Luo, and Lu (2024) Their neglect of nonlocal information, however, leads to systematic errors in key electronic properties: atomic shell structures vanish Wang and Carter (2002), charge density cusps at nuclei are unphysical unless additional correction introduced Kato (1957); Hutcheon and Wibowo-Teale (2024).
Nonlocal KEDFs enhance accuracy by incorporating nonlocal information through convolution kernels $w(\mathbf{r},\mathbf{r}^{\prime})$. The Wang-Teter-like (WT-like) functionals, including WT Wang and Teter (1992), Smargiassi-Madden (SM) Smargiassi and Madden (1994), and Perrot Perrot (1994), employ a density-independent kernel $w(k_{\mathrm{F}}^{0},|\mathbf{r}-\mathbf{r}^{\prime}|)$, parametrized by $k_{\mathrm{F}}^{0}=(3\pi^{2}\rho_{0})^{1/3}$. Here, $\rho_{0}$ is a parameter conventionally set as the average charge density $\rho_{\mathrm{avg}}$. The WT-like KEDFs generally decompose the non-interacting kinetic energy $T_{s}$ as
@@
Distinct physical regimes motivate specific parameterizations: The original WT KEDF ($\alpha$=$\beta$=$\frac{5}{6}$) targets weakly varying uniform electron gas (UEG), while the Perrot KEDF ($\alpha$=$\beta$=1) was constructed for thin electron gas, and the SM KEDF ($\alpha$=$\beta$=$\frac{1}{2}$) is derived from low-$q$ limit. Despite achieving $O(N\ln N)$ scaling and improved accuracy in bulk systems, these functionals catastrophically failure in isolated systems due to the Blanc-Cancès (BC) instability Blanc and Cancès (2005), manifested as total energy unboundedness from below.
Our analysis (see Supplementary Information (SI)) traces the BC instability to the ill-defined nature of $\rho_{\mathrm{avg}}$ as a rigid spatial average that remains fixed under density scaling $\rho_{\sigma}(\mathbf{r})=\sigma^{3}\rho_{1}(\sigma\mathbf{r})$ in isolated systems. This unphysical constant simultaneously violates the exact scaling law $T_{s}[\rho_{\sigma}]=\sigma^{2}T_{s}[\rho_{1}]$ Levy and Perdew (1985) and generates negative Pauli energy $T_{\theta}$ through its mismatch with the physical density $\rho(\mathbf{r})$, contradicting its fundamental positivity requirement Levy and Ou-Yang (1988). The BC instability therefore stems from this dual failure of $\rho_{\mathrm{avg}}$, which both breaks scaling invariance and misrepresents the true density, collectively rendering the functional unbounded. This understanding addresses a longstanding challenge in nonlocal KEDF development.
@@
With the BC instability resolved, we benchmark the ext-WT KEDF against semilocal functionals, including the second-order gradient expansion (GE2) Kirzhnits (1957) and Luo-Karasiev-Trickey (LKT) KEDFs Luo, Karasiev, and Trickey (2018), alongside the nonlocal WT KEDF, and KSDFT for single-atom systems of 56 elements. All calculations were performed within the ABACUS 3.8.0 packages Li et al. (2016); Zhou et al. (2025) with Perdew-Burke-Ernzerhof (PBE) Perdew, Burke, and Ernzerhof (1996) exchange-correclation functional, employing three distinct potentials: bare Coulomb potential for H and helium (He), widely-used bulk-derived local pseudopotentials (BLPS) Huang and Carter (2008) covering nine elements, and recently proposed high-quality local pseudopotentials (HQLPS) Chi and Huang (2024) covering all simple and transition metals.